How do I solve:

FInd a real number, k, such that the line -7x+ky-16=0 has y intercept 13

Answer by lwsshak3(11628)

FInd a real number, k, such that the line -7x+ky-16=0 has y intercept 13
**
equation for a line: y=mx+b, m=slope, b=y-intercept
-7x+ky-16=0
ky=7x+16
y=7x/k+16/k
y-intercept=16/k=13
k=16/13